OpenStudy (anonymous):

A painter used 5/6 gallon of paint for one job and 1/4 gallon for another job. How many gallons of paint did the painter use for both jobs? A. 1/2 gal B. 3/5 gal C. 1 1/12 gal D. 11/6 gal

OpenStudy (anonymous):

You get the answer by adding the two fractions

OpenStudy (anonymous):

not at all the answer is wrong

OpenStudy (anonymous):

none is the answer

OpenStudy (anonymous):

studyisland, why don't you walk us through what you usually do to solve this kind of problem?

OpenStudy (anonymous):

\[\frac{ 5 }{ 6 }+\frac{ 1 }{ 4 }=\frac{ 6 }{ 24 }+\frac{ 6 }{ 24 }=\frac{ 26 }{ }\] Then simplify

OpenStudy (anonymous):

the answer is13/12

OpenStudy (anonymous):

how u simplify the values can u explain

OpenStudy (anonymous):

Whoops i meant to change into a mixed number by simplifying it

OpenStudy (anonymous):

also the answer is \[\frac{ 26 }{ 24 }\]

OpenStudy (anonymous):

by simplifying we get 13/12 so option is correct ok

OpenStudy (anonymous):

You are right BUT

OpenStudy (anonymous):

WE change that to \[1 \frac{ 1 }{ 12 }\]

OpenStudy (anonymous):

which is letter C

OpenStudy (anonymous):

OK?

OpenStudy (anonymous):

c is looking like 11/12

OpenStudy (anonymous):

well its not.

OpenStudy (anonymous):

ok

OpenStudy (cggurumanjunath):